OpenStudy (anonymous):

Solve P = 4(3a – b) for a in terms of P and b. You must show all work.

OpenStudy (anonymous):

P = 4(3a – b) Distribute 4: P = 12a - 4b transfer 4b to the other side P + 4b = 12a -4b + 4b P + 4b = 12a divide both sides by 12 (P+4b)/12 = (12a)/12 (P+4b)/12 = a